OpenStudy (anonymous):

Find the exact sum of the series S = ∑ (5^(n+2))/(2^(n)*pi^(n+1)) for n = 1 to infinity

OpenStudy (campbell_st):

looking at the terms \[T_{1} = \frac{5^3}{2\pi^2}\] \[T_{2} = \frac{5^4}{2^2\pi^3}\] so it would appear the common ratio is \[r = \frac{5}{2\pi}\] since the common ratio is -1 < r < 1 it means the series will have a limiting sum \[S_{\infty} = \frac{a}{1 - r}\] a is the 1st term, shown above and r is the common ratio. Hope this helps.

OpenStudy (anonymous):

thank you!